# Electromagnetic tomography of spin-$\frac{3}{2}$ hidden-charm strange pentaquarks

### v1

@@ I Motivation
The concept of hadronic states beyond the conventional quark-antiquark mesons and three-quark baryons has been proposed for many decades. Experimental verification of such exotic configurations was first achieved in 2003 with the discovery of the $X(3872)$ by the Belle Collaboration [1]. Due to its properties that cannot be explained within the framework of standard mesons, this state is widely regarded as the first confirmed tetraquark candidate. Since then, numerous exotic hadrons featuring nontraditional quark arrangements have been observed, and the variety of these states has steadily expanded. These findings have stimulated extensive theoretical and experimental studies aimed at understanding the nature and internal structure of these unconventional multi-quark systems, which in turn provide valuable information on the dynamics of strong interactions at low energies. A range of theoretical interpretations has been proposed, including compact multiquark models, hadronic molecules, and explanations based on kinematic effects, among others [2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17]. Together, these efforts have significantly advanced our understanding of exotic hadronic matter and continue to guide ongoing research in hadron spectroscopy.
@@ II QCD light-cone sum rules setup
Here, $S^{\rm free}(x)$ corresponds to the leading term of the light and heavy quark propagators.
@@
where $\Gamma_{i}=\mathrm{1},\gamma_{5},\gamma_{\mu},i\gamma_{5}\gamma_{\mu},\sigma_{\mu\nu}/2$. This substitution leads to matrix elements of the form $\langle\gamma(q)|\bar{q}(x)\Gamma_{i}q(0)|0\rangle$ and $\langle\gamma(q)|\bar{q}(x)\Gamma_{i}G_{\mu\nu}q(0)|0\rangle$, which are parameterized in terms of the photon DAs, following the established formalism [63]. By systematically implementing the substitutions from Eqs. (21) and (22), both the perturbative and non-perturbative photon interactions are fully incorporated into the QCD representation of the correlation function. The resulting expressions in coordinate space are then transformed to momentum space via a Fourier transformation, yielding the final form used for the subsequent QCD light-cone sum rule analysis.
@@ III Numerical Illustrations
This section is devoted to the numerical investigation of the QCD light-cone sum rule aimed at evaluating the electromagnetic multipole moments of the $P^{\Lambda}_{\psi s}$ state. In order to extract the magnetic dipole moment, several input parameters must be specified. The set of numerical inputs adopted in this study is summarized in Table 1. For the numerical evaluation, the up and down quark masses are set to zero ($m_{u}=m_{d}=0$), while contributions proportional to $m_{s}$ are retained. The photon DAs, along with their explicit expressions and the necessary numerical parameters, are adopted from Ref. [63]. In the present study, the calculations incorporate photon DAs up to twist-4 in order to account consistently for higher-twist effects.
@@
   $\langle\bar{s}s\rangle$    $0.8\times(-0.24\pm 0.01)^{3}$ GeV3 [65]
@@
Table 3: Numerical results of the magnetic dipole (in units of $\mu_{N}$), electric quadrupole (in units of $10^{-2}$ fm2) and magnetic octupole (in units of $10^{-2}$ fm3) moments of the $P^{\Lambda}_{\psi s}$.
State $\mu_{P^{\Lambda}_{\psi s}}$ $\mathcal{Q}_{P^{\Lambda}_{\psi s}}$ $\mathcal{O}_{P^{\Lambda}_{\psi s}}$

### v3

@@ I Motivation
The concept of hadronic states beyond the conventional quark-antiquark mesons and three-quark baryons has been proposed for many decades. Experimental verification of such exotic configurations was first achieved in 2003 with the discovery of the $X(3872)$ by the Belle Collaboration [22]. Due to its properties that cannot be explained within the framework of standard mesons, this state is widely regarded as the first confirmed tetraquark candidate. Since then, numerous exotic hadrons featuring nontraditional quark arrangements have been observed, and the variety of these states has steadily expanded. These findings have stimulated extensive theoretical and experimental studies aimed at understanding the nature and internal structure of these unconventional multi-quark systems, which in turn provide valuable information on the dynamics of strong interactions at low energies. A range of theoretical interpretations has been proposed, including compact multiquark models, hadronic molecules, and explanations based on kinematic effects, among others [27, 26, 45, 33, 43, 14, 6, 20, 8, 30, 37, 74, 24, 25, 19, 38, 7, 15]. Together, these efforts have significantly advanced our understanding of exotic hadronic matter and continue to guide ongoing research in hadron spectroscopy.
@@ II QCD light-cone sum rules setup
Here, $S^{\rm free}(x)$ corresponds to the leading term of the light- and heavy quark propagators.
@@
where $\Gamma_{i}=\mathrm{1},\gamma_{5},\gamma_{\mu},i\gamma_{5}\gamma_{\mu},\sigma_{\mu\nu}/2$. This substitution leads to matrix elements of the form $\langle\gamma(q)|\bar{q}(x)\Gamma_{i}q(0)|0\rangle$ and $\langle\gamma(q)|\bar{q}(x)\Gamma_{i}G_{\mu\nu}q(0)|0\rangle$, which are parameterized in terms of the DAs of the photon, following the established formalism [12]. By systematically implementing the substitutions from Eqs. (21) and (22), both the perturbative and non-perturbative photon interactions are fully incorporated into the QCD representation of the correlation function. The resulting expressions in coordinate space are then transformed to momentum space via a Fourier transformation, yielding the final form used for the subsequent QCD light-cone sum rule analysis.
@@ III Numerical Illustrations
This section is devoted to the numerical investigation of the QCD light-cone sum rule aimed at evaluating the electromagnetic multipole moments of the $P^{\Lambda}_{\psi s}$ state. In order to extract the magnetic dipole moment, several input parameters are required. These include the pentaquark masses ($m_{J^{i}_{\mu}}$) and residues ($\lambda_{J^{i}_{\mu}}$), the quark masses ($m_{u},m_{d},m_{s},m_{c}$), the quark condensates ($\langle\bar{q}q\rangle$ and $\langle\bar{s}s\rangle$), the gluon condensate ($\langle g_{s}^{2}G^{2}\rangle$), the magnetic susceptibility of the quark condensate $\chi$, and $f_{3\gamma}$—a non-perturbative constant measuring the chirality-flipping soft part of a photon’s structure. The set of numerical inputs adopted in this study is summarized in Table 1. For the numerical evaluation, the up and down quark masses are set to zero ($m_{u}=m_{d}=0$), while contributions proportional to $m_{s}$ are retained. The photon DAs, along with their explicit expressions and the necessary numerical parameters, are adopted from [12]. In the present study, the calculations incorporate photon DAs up to twist-4 in order to account consistently for higher-twist effects.
@@
   $\langle\bar{s}s\rangle$    $0.8\times(-0.24\pm 0.01)^{3}$~~ GeV3 [31]
@@
Table 3: Numerical results of the magnetic dipole (in units of $\mu_{N}$), electric quadrupole (in units of fm2) and magnetic octupole (in units of fm3) moments of the $P^{\Lambda}_{\psi s}$.
State $\mu_{P^{\Lambda}_{\psi s}}$ $\mathcal{Q}_{P^{\Lambda}_{\psi s}}(\times 10^{-2})$ $\mathcal{O}_{P^{\Lambda}_{\psi s}}(\times 10^{-2})$
@@ V Summary and Conclusions
  • Future studies employing currents explicitly constructed for molecular configurations (e.g., $\bar{D}^{(*)}\Xi_{c}$ meson-baryon pairs) would provide a direct comparison with the present compact-diquark based results, thereby offering a more definitive test for the internal structure of the $P^{\Lambda}_{\psi s}$ states.
@@
The present framework can be systematically extended to investigate the electromagnetic properties of non-strange hidden-charm pentaquarks, as well as those with $J^{P}=\frac{1}{2}^{-}$, by constructing appropriate interpolating currents and repeating the analysis outlined in this work. Such studies would provide a more complete spectroscopic map and further test the structural sensitivity of multipole moments.